ghostmasterTy:

1. 2 2/7 + 2 1/3 2. 2 1/8 +3 1/4 3. 6 1/7 +3 1/14 4. 1 3/8 + 4 3/16 5. 10 1/9 3 2/27 6. 5 3/8 + 1 2/64 7. 4 4/5 4 1/6 8. 5 2/5 + 5 1/4 9. 9 1/3 + 5 2/5 10. 7 3/4 + 2 3/5

ghostmasterTy:

im not trying to spam or get levels

itsmehjay:

Are these supposed to be mixed numbers?

ghostmasterTy:

i need help with my homework that was the old me

ghostmasterTy:

but can i be crispycattho

crispyrat:

so this is the same idea but e convert them for the same denominater for example 1/4*2/2=2/8 2 1/8+3 1/4=2 1/8+3 2/8=5 5/8

ghostmasterTy:

@itsmehjay wrote:
Are these supposed to be mixed numbers?
Adding fractions

itsmehjay:

1. Find the same denominater 2. add the numerators once they have the same bottom 3. also add the whole numbers 4. and boom-

itsmehjay:

so lets do one together

ghostmasterTy:

ok and crispy rat can help

ghostmasterTy:

to

ghostmasterTy:

2/32 = 435645/656 right

itsmehjay:

the first one is- 2 2/7 + 2 1/3 2/7 and 1/3 don't have the same denominater so you cant add them both

ghostmasterTy:

ok

ghostmasterTy:

soooooooooooo

itsmehjay:

you need to find a common denominater for the both of them

ghostmasterTy:

its 2

itsmehjay:

what number fits both for 7 and 3?

itsmehjay:

@ghostmasterty wrote:
its 2
how did you get that?

itsmehjay:

3 doesn't go into 7 so lets find a common multiple

itsmehjay:

what is the least common multiple of 3 and 7? @ghostmasterty

ghostmasterTy:

3

ghostmasterTy:

ghostmasterTy:

ghostmasterTy:

pl shelp me

ghostmasterTy:

you help

ghostmasterTy:

im done you at helping closeing it go to the next one

ghostmasterTy:

ghostmasterTy:

ghostmasterTy:

itsmehjay:

Stop it-

ghostmasterTy:

ok

Camm:

XD